# An optimal parallel-in-time preconditioner for parabolic optimal control problems

### v1

@@ Abstract
In this work, we propose a novel diagonalization-based preconditioner for the all-at-once linear system arising from the optimal control problem of parabolic equations. The proposed preconditioner is constructed based on an $\epsilon$-circulant modification to the rotated block diagonal (RBD) preconditioning technique, which can be efficiently diagonalized by fast Fourier transforms in a parallel-in-time fashion. To our knowledge, this marks the first application of the $\epsilon$-circulant modification to RBD preconditioning. Before our work, the studies of PinT preconditioning techniques for the optimal control problem are mainly focused on $\epsilon$-circulant modification to Schur complement based preconditioners, which involves multiplication of forward and backward evolutionary processes and thus square the condition number. Compared with those Schur complement based preconditioning techniques in the literature, the advantage of the proposed $\epsilon$-circulant modified RBD preconditioning is that it does not involve the multiplication of forward and backward evolutionary processes. When the generalized minimal residual method is deployed on the preconditioned system, we prove that when choosing $\epsilon=\mathcal{O}(\sqrt{\tau})$ with $\tau$ being the temporal step-size , the convergence rate of the preconditioned GMRES solver is independent of the matrix size and the regularization parameter. Such restriction on $\epsilon$ is more relax than the assumptions on $\epsilon$ from other works related to $\epsilon$-circulant based preconditioning techniques for the optimal control problem. Numerical results are provided to demonstrate the effectiveness of our proposed solvers.
@@
Block Toeplitz matrices, parallel-in-time, $\epsilon$-circulant matrices, preconditioning, all-at-once systems
@@ 1 Introduction
Developing diagonalization-based parallel-in-time (PinT) for solving optimization problems constrained by partial differential equations (PDEs) has gained substantial attention in recently years. We refer to [19, 12, 27, 9] and the references therein for a comprehensive overview of these optimization problems. Various efficient PinT preconditioners have been proposed to solve all-at-once linear systems arising from time-dependent PDEs [23, 22, 17, 20, 15] and PDE-constrained optimal control problems [14, 29, 30, 31, 13, 8, 10]; see for a review paper on these diagonalization-based methods [21] and the references therein reported.
@@
where $u,g\in L^{2}$ are the distributed control and the targeted tracking trajectory, respectively, $\gamma>0$ is a regularization parameter, $\mathcal{L}=\nabla\cdot(a({x})\nabla)$, and both $f$ and $y_{0}$ are given functions. The theoretical aspects of existence, uniqueness, and regularity of the solution, under suitable assumptions, were thoroughly studied in [19]. The optimal solution of (1) & (2) can be characterized by the following system:
(3) $\left\{\begin{array}[]{lc}y_{t}-\mathcal{L}y-\frac{1}{\gamma}p=f,\quad(x,t)\in\Omega\times(0,T],\qquad y=0,\quad(x,t)\in\partial\Omega\times(0,T],\\ y(x,0)=y_{0}\quad x\in\Omega,\\ -p_{t}-\mathcal{L}p+y=g,\quad(x,t)\in\Omega\times(0,T],\qquad p=0,\quad(x,t)\in\partial\Omega\times(0,T],\\ p(x,T)=0\quad x\in\Omega,\end{array}\right.\,$
where the control variable $u$ has been eliminated.
Similar to [31, 30], we discretize (3) using the backward Euler method for time and a typical finite element method for space discretization, which gives
$M_{m}\frac{\mathbf{y}_{m}^{(k+1)}-\mathbf{y}_{m}^{(k)}}{\tau}+K_{m}\mathbf{y}_{m}^{(k+1)}$ $=$ $M_{m}\left(\mathbf{f}_{m}^{(k+1)}+\frac{1}{\gamma}\mathbf{p}_{m}^{(k)}\right),$
$-M_{m}\frac{\mathbf{p}_{m}^{(k+1)}-\mathbf{p}_{m}^{(k)}}{\tau}+K_{m}\mathbf{p}_{m}^{(k)}$ $=$ $M_{m}\left(\mathbf{g}_{m}^{(k)}-\mathbf{y}_{m}^{(k+1)}\right).$
Combining the given initial and boundary conditions, one needs to solve the following linear system
@@
where we have $\mathbf{y}=[\mathbf{y}_{m}^{(1)},\cdots,\mathbf{y}_{m}^{(n)}]^{\top}$, $\mathbf{p}=[\mathbf{p}_{m}^{(0)},\cdots,\mathbf{p}_{m}^{(n-1)}]^{\top}$,
$\mathbf{f}=\begin{bmatrix}M_{m}\tau\mathbf{f}_{m}^{(1)}+M_{m}\mathbf{y}_{m}^{(0)}\\ M_{m}\tau\mathbf{f}_{m}^{(2)}\\ \vdots\\ M_{m}\tau\mathbf{f}_{m}^{(n)}\\ \end{bmatrix},\quad\mathbf{g}=\tau\begin{bmatrix}M_{m}\mathbf{g}_{m}^{(0)}\\ M_{m}\mathbf{g}_{m}^{(1)}\\ \vdots\\ M_{m}\mathbf{g}_{m}^{(n-1)}\\ \end{bmatrix},$
(5) $\widetilde{\mathcal{A}}$ $=$ $\begin{bmatrix}\tau I_{n}\otimes M_{m}&B_{n}^{\top}\otimes M_{m}+\tau I_{n}\otimes K_{m}\\ B_{n}\otimes M_{m}+\tau I_{n}\otimes K_{m}&-\frac{\tau}{\gamma}I_{n}\otimes M_{m}\end{bmatrix},$
and the matrix $B_{n}\in\mathbb{R}^{n\times n}$ is as follows.
(6) $B_{n}=\begin{bmatrix}1&&&&\\ -1&1&&&\\ &-1&1&&\\ &&\ddots&\ddots&\\ &&&-1&1\end{bmatrix}.$
The matrices $M_{m}$ and $K_{m}$ represent the mass matrix and the stiffness matrix, respectively. For the finite difference method, we have $M_{m}=I_{m}$ and $K_{m}=-L_{m}$, where $-L_{m}$ is the discretization matrix of the negative Laplacian. Throughout, the following conditions on both $M_{m}$ and $K_{m}$ are assumed, which are in line with [17, Section 2]:
@@
We then further transform (5) into the following equivalent system
(8) ${\mathcal{A}}\underbrace{\begin{bmatrix}\sqrt{\gamma}\mathbf{y}\\ \mathbf{p}\end{bmatrix}}_{:={\bf x}}=\underbrace{\begin{bmatrix}\mathbf{g}\\ -\sqrt{\gamma}\mathbf{f}\end{bmatrix}}_{:={\bf b}},$
@@
In what follows, we will develop a preconditioned generalized minimal residual (GMRES) method for the nonsymmetric system (8).
For $\mathcal{A}$, we first consider the following rotated block diagonal (RBD) preconditioner proposed in [4]:
@@
where $\mathcal{T}$ is defined in (10). As will be shown in Theorem 2.4, the matrix $\mathcal{P}$ can achieve an excellent preconditioning effect for $\mathcal{A}$. Yet, $\mathcal{P}$ is in general computationally expensive to invert. Thus, we propose the following PinT preconditioner (or denoted by RBD $\epsilon$-circulant preconditioner) as a modification of $\mathcal{P}$, which can be efficiently implemented:
@@
$D_{\epsilon}={\rm diag}\left(\epsilon^{\frac{i-1}{N}}\right)_{i=1}^{n},$
@@
Several efficient preconditioned iterative solvers have been proposed for (5). One notable class of such methods is the circulant-based preconditioners [31, 8, 10], whose main idea is to replace the Toeplitz matrix $B_{n}$ defined by (6) by a circulant or skew-circulant matrix in order to form an efficient preconditioner. Despite their numerically observed success, theoretically these preconditioners have not shown to be parameter-robust due to the low-rank approximation from the circulant-based approximation.
@@
Another major existing method is preconditioning technique based on Schur complement approximations [30, 14, 18], which is a typical preconditioning approach in the context of preconditioning for saddle point systems (see, e.g., [26, 28]) and their effectiveness is based on approximating the Schur complements by multiplications of easily invertible matrices (e.g., [25, 14, 2, 24]). Recently, $\epsilon$-circulant modifications have been introduced to such approximate-Schur-complement preconditioning techniques for the optimal control problems, which leads to a significant improvement on computational efficiency and parallelism along time dimension (see, e.g, [17, 20, 15, 16, 30, 21]).
@@
Our proposed preconditioning method, while still based on $\epsilon$-circulant matrices, diverges from the approximate-Schur-complement approach. Specifically, the Toeplitz matrix $B_{n}$ is formulated using an $\epsilon$-circulant matrix as detailed in (13). An important advancement of our method is that it can achieve optimal convergence rate when $\epsilon=\mathcal{O}(\tau^{1/2})$. This contrasts with $\epsilon$-circulant modified-approximate-Schur-complement preconditioner, such as described in [18], which requires a more restrictive $\epsilon=\mathcal{O}(\tau^{2})$. Additionally, while the other $\epsilon$-circulant modified-approximate-Schur-complement approach in [30] also allows for $\epsilon=\mathcal{O}(\tau^{1/2})$, it imposes the severe theoretical constraint that $n$ must be an even number, a limitation our method does not share.
@@
In addition to the RBD preconditioner of interest in this work, other effective preconditioners such as the preconditioned square block matrix [1] could potentially be incorporated with the $\epsilon$-circulant preconditioning technique. We have chosen the RBD preconditioner due to its simple form, whose inversion only mainly determined by that of the block diagonal with $\mathcal{C}_{\epsilon}+\alpha{I}_{n}\otimes M_{m}$ and $\mathcal{C}^{\top}_{\epsilon}+\alpha{I}_{n}\otimes M_{m}$. Such an inversion can be straightforwardly implemented in a PinT way.
@@ 2 Main Result
To check the invertibility of $\mathcal{P}_{\epsilon}$, it suffices to verify the invertibility of $\mathcal{G}$ and $\mathcal{H}_{\epsilon}$ respectively. Clearly, $\mathcal{G}$ is invertible. By (14), it is straightforward to see that $\mathcal{C}_{\epsilon}+\alpha I_{n}\otimes M_{m}$ is similar to a block diagonal matrix with each diagonal block having the form $\alpha M_{m}+z_{1}I_{m}+z_{2}{\bf i}I_{m}$ ($z_{1}\geq 0$, $z_{2}\in\mathbb{R}$). Consequently, $\mathcal{C}_{\epsilon}+\alpha I_{n}\otimes M_{m}$ is invertible. Moreover, $\mathcal{C}_{\epsilon}^{\top}+\alpha{I}_{n}\otimes M_{m}$ as the transpose of $\mathcal{C}_{\epsilon}+\alpha{I}_{n}\otimes M_{m}$ is also invertible. Thus, $\mathcal{H}_{\epsilon}$, as a block diagonal matrix with $\mathcal{C}_{\epsilon}^{\top}+\alpha{I}_{n}\otimes M_{m}$ and $\mathcal{C}_{\epsilon}+\alpha{I}_{n}\otimes M_{m}$ as its diagonal blocks, is also invertible. Therefore, $\mathcal{P}_{\epsilon}$ is invertible.
@@
The convergence behavior of GMRES solver is closely related to the Krylov subspace. For a square matrix $\mathbf{E}\in\mathbb{R}^{k\times k}$ and a vector ${\bf y}\in\mathbb{R}^{k\times 1}$, a Krylov subspace of degree $j\geq 1$ is defined as follows
@@
For a non-singular $k\times k$ real linear system ${\bf Z}{\bf y}={\bf w}$, let ${\bf y}_{j}$ be the GMRES iterative solution at the $j$-th iteration $(j\geq 1)$ with ${\bf y}_{0}$ as an initial guess. Then, the $j$-th iteration solution ${\bf y}_{j}$ minimize the residual error over the Krylov subspace $\mathcal{K}_{j}({\bf Z},{\bf r}_{0})$ with ${\bf r}_{0}={\bf w}-{\bf Z}{\bf y}_{0},$ i.e.,
@@
Let $\tilde{\mathbf{x}}_{0}$ be arbitrary initial guess for (16). Let $\mathbf{x}_{0}:=\mathcal{G}^{-1}\mathcal{W}^{-\frac{1}{2}}\tilde{\mathbf{x}}_{0}$ be the initial guess for (15). Let $\mathbf{x}_{j}(\tilde{\mathbf{x}}_{j},\textnormal{respectively})$ be the j-th $(j\geq 1)$ iteration solution derived by applying GMRES solver to (15) ((16),respectively)italic-(16italic-)$,\textnormal{respectively})$ ) with $\mathbf{u}_{0}(\tilde{\mathbf{u}}_{0},\textnormal{respectively})$ as an initial guess. Then,
$\left\|\mathbf{r}_{j}\right\|_{2}\leq\sqrt{2}||\mathcal{W}^{-\frac{1}{2}}||_{2}||\tilde{\bf r}_{j}||_{2},\quad j=1,2,...,$
@@
By definition of $\tilde{\bf x}_{j}$ and $\tilde{\bf x}_{0}$, we know that $\tilde{\bf x}_{j}-\tilde{\bf x}_{0}\in\mathcal{K}_{j}(\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}},\tilde{\mathbf{r}}_{0})$ with $\tilde{\mathbf{r}}_{0}=\tilde{\bf b}-\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}}\tilde{\mathbf{x}}_{0}$, meaning that there exists real numbers $\eta_{0},\eta_{1},...,\eta_{j-1}$ such that
@@
Denote by $\sigma({\bf C})$, the spectrum of a square matrix ${\bf C}$.
@@
Considering ~$\widetilde{\mathcal{E}}=(\widetilde{\mathcal{T}}+\alpha I_{mn})^{-1}(-\widetilde{\mathcal{T}}+\alpha I_{mn})$ with its singular value decomposition, namely, ~$\widetilde{\mathcal{E}}=\mathcal{U}\Sigma\mathcal{V}^{*}$. Then, we can further decompose ~$\widetilde{\mathcal{H}}^{-1}\widetilde{\mathcal{B}}$ as follows:
~$\widetilde{\mathcal{H}}^{-1}\widetilde{\mathcal{B}}$ $=$ $\begin{bmatrix}I_{mn}&-\widetilde{\mathcal{E}}^{\top}\\ \widetilde{\mathcal{E}}&I_{mn}\end{bmatrix}=\begin{bmatrix}\mathcal{V}&\\ &\mathcal{U}\end{bmatrix}\begin{bmatrix}I_{mn}&-\Sigma^{*}\\ \Sigma&I_{mn}\end{bmatrix}\begin{bmatrix}\mathcal{V}^{*}&\\ &\mathcal{U}^{*}\end{bmatrix}.$

### v2

# An optimal diagonalization-based preconditioner for parabolic optimal control problems

@@ Abstract
In this work, we propose a novel diagonalization-based preconditioner for the all-at-once linear system arising from the optimal control problem of parabolic equations. The proposed preconditioner is constructed based on an $\epsilon$-circulant modification to the rotated block diagonal (RBD) preconditioning technique and can be efficiently diagonalized by fast Fourier transforms in a parallel-in-time fashion. To our knowledge, this marks the first application of the $\epsilon$-circulant modification to RBD preconditioning. Before our work, the studies of parallel-in-time preconditioning techniques for the optimal control problem are mainly focused on $\epsilon$-circulant modification to Schur complement based preconditioners, which involves multiplication of forward and backward evolutionary processes and thus square the condition number. Compared with those Schur complement based preconditioning techniques in the literature, the advantage of the proposed $\epsilon$-circulant modified RBD preconditioning is that it does not involve the multiplication of forward and backward evolutionary processes. When the generalized minimal residual method is deployed on the preconditioned system, we prove that when choosing $\epsilon=\mathcal{O}(\sqrt{\tau})$ with $\tau$ being the temporal step-size , the convergence rate of the preconditioned GMRES solver is independent of the matrix size and the regularization parameter. Numerical results are provided to demonstrate the effectiveness of our proposed solvers.
@@
Block Toeplitz matrices, fast Fourier transforms, $\epsilon$-circulant matrices, preconditioning, rotated block diagonal preconditioners
@@ 1 Introduction
Developing diagonalization-based parallel-in-time (PinT) methods for solving optimization problems constrained by partial differential equations (PDEs) has gained substantial attention in recently years. We refer to [22, 15, 32, 10] and the references therein for a comprehensive overview of these optimization problems. Various efficient PinT preconditioners have been proposed to solve all-at-once linear systems arising from time-dependent PDEs [26, 25, 20, 23, 18, 11] and PDE-constrained optimal control problems [17, 34, 35, 36, 16, 8, 13]; for a review paper on these diagonalization-based methods, see [24] and the references therein reported.
@@
where $u,g\in{\color[rgb]{0,0,0}L^{2}(\Omega\times(0,T))}$ are the distributed control and the targeted tracking trajectory, respectively, $\gamma>0$ is a regularization parameter, $\mathcal{L}=\nabla\cdot(a({x})\nabla)$, and both $f$ and $y_{0}$ are given functions. The theoretical aspects of existence, uniqueness, and regularity of the solution, under suitable assumptions, were thoroughly studied in [22]. Following a standard argument [22], the optimal solution of (1) - (2) can be characterized by the following system:
(3) $\left\{\begin{array}[]{lc}y_{t}-\mathcal{L}y-\frac{1}{\gamma}p=f,\quad(x,t)\in\Omega\times(0,T],\qquad y=0,\quad(x,t)\in\partial\Omega\times(0,T],\\ y(x,0)=y_{0},\quad x\in\Omega,\\ -p_{t}-\mathcal{L}p+y=g,\quad(x,t)\in\Omega\times(0,T],\qquad p=0,\quad(x,t)\in\partial\Omega\times(0,T],\\ p(x,T)=0,\quad x\in\Omega,\end{array}\right.\,$
where $p$ denotes the Lagrange multiplier, also known as the co-state variable. In deriving the second PDE in (3), the control variable $u$ is eliminated using the optimality condition $\gamma u-p=0$. For further details, we refer the reader to [35, Section 3.1] and [22].
Following [17], we discretize (3) using the backward Euler method in time with a uniform step size $\tau=T/N$, and apply some spatial discretization with mesh size parameter $m$, which gives
$M_{m}\frac{\mathbf{y}_{m}^{(k_{1}+1)}-\mathbf{y}_{m}^{(k_{1})}}{\tau}+K_{m}\mathbf{y}_{m}^{(k_{1}+1)}$ $=$ $M_{m}\left(\mathbf{f}_{m}^{(k_{1}+1)}+\frac{1}{\gamma}\mathbf{p}_{m}^{(k_{1}+1)}\right),\qquad k_{1}=0,1,2,\dots,N-1,$
$-M_{m}\frac{\mathbf{p}_{m}^{(k_{2}+1)}-\mathbf{p}_{m}^{(k_{2})}}{\tau}+K_{m}\mathbf{p}_{m}^{(k_{2})}$ $=$ $M_{m}\left(\mathbf{g}_{m}^{(k_{2})}-\mathbf{y}_{m}^{(k_{2})}\right),\qquad k_{2}=0,1,2,\dots N-1,$
where the matrices $M_{m}$ and $K_{m}$ represent the mass matrix and the stiffness matrix, respectively, if a finite element method is used. For the finite difference method, we have $M_{m}=I_{m}$ and $K_{m}=-L_{m}$, where $-L_{m}$ is the discretization matrix of the negative Laplacian.
Combining the given initial and final conditions that $\mathbf{y}_{m}^{(0)}=y_{0}$ and $\mathbf{p}_{m}^{(N)}=\mathbf{0}$, one needs to solve the following all-at-once system assembled using the formulas above for $k_{1}=0,1,\dots,N-2$ and $k_{2}=1,2,\dots,N-1$ (see, e.g., [8, 36], where the system is derived in the same way):
@@
where we have $\mathbf{y}=[\mathbf{y}_{m}^{(1)},\cdots,\mathbf{y}_{m}^{(N-1)}]^{\top}$, $\mathbf{p}=[\mathbf{p}_{m}^{(1)},\cdots,\mathbf{p}_{m}^{(N-1)}]^{\top}$,
$\mathbf{f}=\begin{bmatrix}\tau M_{m}\mathbf{f}_{m}^{(1)}+M_{m}\mathbf{y}_{m}^{(0)}\\ \tau M_{m}\mathbf{f}_{m}^{(2)}\\ \vdots\\ \tau M_{m}\mathbf{f}_{m}^{(N-1)}\\ \end{bmatrix},\quad\mathbf{g}=\begin{bmatrix}\tau M_{m}\mathbf{g}_{m}^{(1)}\\ \vdots\\ \tau M_{m}\mathbf{g}_{m}^{(N-2)}\\ \tau M_{m}\mathbf{g}_{m}^{(N-1)}+M_{m}\mathbf{p}_{m}^{(N)}\\ \end{bmatrix},$
(5) $\widetilde{\mathcal{A}}$ $=$ $\begin{bmatrix}\tau I_{N-1}\otimes M_{m}&B_{N-1}^{\top}\otimes M_{m}+\tau I_{N-1}\otimes K_{m}\\ B_{N-1}\otimes M_{m}+\tau I_{N-1}\otimes K_{m}&-\frac{\tau}{\gamma}I_{N-1}\otimes M_{m}\end{bmatrix},$
and the matrix $B_{N-1}\in\mathbb{R}^{N-1\times N-1}$ is as follows.
(6) $B_{N-1}=\begin{bmatrix}1&&&&\\ -1&1&&&\\ &-1&1&&\\ &&\ddots&\ddots&\\ &&&-1&1\end{bmatrix}.$
After solving the system (4) for $\mathbf{y}$ and $\mathbf{p}$, one proceeds to compute $\mathbf{y}_{m}^{(N)}$ and $\mathbf{p}_{m}^{(0)}$ via the following discretization formulas for $k_{1}=N-1$ and $k_{2}=0$:
$(M_{m}+\tau K_{m})\mathbf{y}_{m}^{(N)}$ $=$ $M_{m}\mathbf{y}_{m}^{(N-1)}+\tau M_{m}\mathbf{f}_{m}^{(N)}+\frac{\tau}{\gamma}M_{m}\mathbf{p}_{m}^{(N)},$
$(M_{m}+\tau K_{m})\mathbf{p}_{m}^{(0)}$ $=$ $M_{m}\mathbf{p}_{m}^{(1)}+\tau M_{m}\mathbf{g}_{m}^{(0)}-{\tau}M_{m}\mathbf{y}_{m}^{(0)}.$
The two linear systems above are classical elliptic problems, which can be efficiently solved with linear complexity by solvers proposed in the literature, see, e.g., [9, 12] and the references therein.
Throughout, the following conditions on both $M_{m}$ and $K_{m}$ are assumed, which are in line with [20, Section 2]:
@@
By letting $n=N-1$, we further transform (5) into the following equivalent system , for which our proposed preconditioner is specifically developed:
(8) ${\mathcal{A}}\underbrace{\begin{bmatrix}\sqrt{\gamma}\mathbf{y}\\ \mathbf{p}\end{bmatrix}}_{{\color[rgb]{0,0,0}=:}{\bf x}}=\underbrace{\begin{bmatrix}\mathbf{g}\\ -\sqrt{\gamma}\mathbf{f}\end{bmatrix}}_{{\color[rgb]{0,0,0}=:}{\bf b}},$
@@
In what follows, we will develop a preconditioner for the nonsymmetric system (8), to be applied within GMRES [31].
For $\mathcal{A}$, we first consider the following rotated block diagonal (RBD) preconditioner firstly proposed in [4]:
@@
where $\mathcal{T}$ is defined in (10). As will be shown in Theorem 2.4, the matrix $\mathcal{P}$ can achieve an excellent preconditioning effect for $\mathcal{A}$. Yet, $\mathcal{P}$ is in general computationally expensive to invert in a PinT manner. Thus, we propose the following PinT preconditioner (denoted by RBD $\epsilon$-circulant preconditioner) as a modification of $\mathcal{P}$, which can be efficiently implemented in parallel:
@@
$D_{\epsilon}={\rm diag}\left(\epsilon^{\frac{i-1}{{\color[rgb]{0,0,0}n}}}\right)_{i=1}^{n},$
@@
Several efficient preconditioned iterative solvers have been proposed for (5). One notable class of such methods is the circulant-based preconditioners [36, 8, 13], whose main idea is to replace the Toeplitz matrix $B_{n}$ defined by (6) by a circulant or skew-circulant matrix in order to form an efficient preconditioner. Despite their numerically observed success, theoretically these preconditioners have not shown to be parameter-robust in theory, meaning that their performance can degrade significantly as problem parameters—such as the mesh size or the regularization parameter, primarily due to the low-rank term in the circulant-based approximation.
@@
Another major existing method is preconditioning technique based on Schur complement approximations [29, 35, 17, 21], which is a typical preconditioning approach in the context of preconditioning for saddle point systems (see, e.g., [27, 33]) and their effectiveness is based on approximating the Schur complements by multiplications of easily invertible matrices (e.g., [28, 17, 2]). Recently, $\epsilon$-circulant modifications have been introduced to such approximate-Schur-complement preconditioning techniques for the optimal control problems, which leads to a significant improvement on computational efficiency and parallelism along the time dimension (see, e.g, [20, 23, 18, 19, 35, 24]).
@@
Our proposed preconditioning method, while still based on $\epsilon$-circulant matrices, diverges from the approximate-Schur-complement approach. Specifically, the Toeplitz matrix $B_{n}$ is replaced with an $\epsilon$-circulant matrix, as detailed in (13). A key advancement of our method is that it achieves an optimal convergence rate when $\epsilon=\mathcal{O}(\sqrt{\tau})$. This sets it apart from the $\epsilon$-circulant modified-approximate-Schur-complement preconditioners proposed in [21, 35].
We note that both [21, 35] focus on the second-order Crank–Nicolson scheme, while our work concerns the backward Euler scheme. On one hand, it remains unclear whether their methods can be directly applied to our setting. On the other hand, although the time discretizations differ, the underlying challenge of constructing effective $\epsilon$-circulant-based preconditioners is closely related. In particular, the choice of $\epsilon$ plays a critical role in balancing approximation quality and numerical stability in both contexts. Therefore, comparing the required conditions on $\epsilon$ remains meaningful. Specifically, the method in [21] requires a more restrictive condition $\epsilon=\mathcal{O}(\tau^{2})$. While the approach in [35] also allows $\epsilon=\mathcal{O}(\sqrt{\tau})$, it imposes a strong theoretical constraint that $n$ must be even—a limitation that our method does not impose.
Furthermore, it is worth noting that our preconditioning approach differs from that in [17], even though both resulting linear systems are derived using the same discretization scheme. Specifically, by enforcing initial and final time conditions as $(M_{m}+\tau K_{m})\mathbf{y}_{m}^{(0)}=(M_{m}+\tau K_{m}){y}_{0}$ and $(M_{m}+\tau K_{m})\mathbf{y}_{m}^{(N)}=\mathbf{0}$, the system in [17] is formulated as:
(15) $\begin{bmatrix}\tau\hat{I}_{N+1}\otimes M_{m}&B_{N+1}^{\top}\otimes M_{m}+\tau I_{N+1}\otimes K_{m}\\ B_{N+1}\otimes M_{m}+\tau I_{N+1}\otimes K_{m}&-\frac{\Delta t}{\gamma}\check{I}_{N+1}\otimes M_{m}\\ \end{bmatrix}\begin{bmatrix}\boldsymbol{y}\\ \boldsymbol{p}\\ \end{bmatrix}=\begin{bmatrix}\boldsymbol{g}\\ \boldsymbol{f}\\ \end{bmatrix},$
with $\boldsymbol{y}=[\mathbf{y}_{m}^{(0)},\mathbf{y}_{m}^{(1)},\cdots,\mathbf{y}_{m}^{(N)}]^{\top}$, $\boldsymbol{p}=[\mathbf{p}_{m}^{(0)},\cdots,\mathbf{p}_{m}^{(N)}]^{\top}$,
$\boldsymbol{f}=\begin{bmatrix}(M_{m}+\tau K_{m})y_{0}\\ \tau M_{m}\mathbf{f}_{m}^{(1)}\\ \vdots\\ \tau M_{m}\mathbf{f}_{m}^{(N-1)}\\ \end{bmatrix},\quad\boldsymbol{g}=\begin{bmatrix}\tau M_{m}\mathbf{g}_{m}^{(0)}\\ \vdots\\ \tau M_{m}\mathbf{g}_{m}^{(N-1)}\\ \mathbf{0}\\ \end{bmatrix},$
and
$\hat{I}_{N+1}=\begin{bmatrix}1&&&\\ &\ddots&&\\ &&1&\\ &&&0\\ \end{bmatrix}\in\mathbb{R}^{N+1\times N+1},\quad\quad\check{I}_{N+1}=\begin{bmatrix}0&&&\\ &1&&\\ &&\ddots&\\ &&&1\\ \end{bmatrix}\in\mathbb{R}^{N+1\times N+1}.$
A Schur-complement-based preconditioner was then proposed in [17] for the system (15); however, its parallelization was not discussed. In contrast, our proposed preconditioner for the system (4) can be implemented efficiently using fast Fourier transforms and is well-suited for parallel execution.
@@
In addition to the RBD preconditioner of interest in this work, other effective preconditioners such as the preconditioned square block matrix [1] could potentially be incorporated with the $\epsilon$-circulant preconditioning technique. We have chosen the RBD preconditioner due to its simple form, whose inversion is only mainly determined by that of the block diagonal with $\mathcal{C}_{\epsilon}+\alpha{I}_{n}\otimes M_{m}$ and $\mathcal{C}^{\top}_{\epsilon}+\alpha{I}_{n}\otimes M_{m}$. Such an inversion can be straightforwardly implemented in a PinT way.
@@ 2 Main Result
To check the invertibility of $\mathcal{P}_{\epsilon}$, it suffices to verify the invertibility of $\mathcal{G}$ and $\mathcal{H}_{\epsilon}$ respectively. Clearly, $\mathcal{G}$ is invertible. By (14), it is straightforward to see that $\mathcal{C}_{\epsilon}+\alpha I_{n}\otimes M_{m}$ is similar to a block diagonal matrix with each diagonal block having the form $\alpha M_{m}+{\color[rgb]{0,0,0}\tau K_{m}}+z_{1}I_{m}+z_{2}{\bf i}I_{m}$ ($z_{1}\geq 0$, $z_{2}\in\mathbb{R}$). Consequently, $\mathcal{C}_{\epsilon}+\alpha I_{n}\otimes M_{m}$ is invertible. Moreover, $\mathcal{C}_{\epsilon}^{\top}+\alpha{I}_{n}\otimes M_{m}$ as the transpose of $\mathcal{C}_{\epsilon}+\alpha{I}_{n}\otimes M_{m}$ is also invertible. Thus, $\mathcal{H}_{\epsilon}$, as a block diagonal matrix with $\mathcal{C}_{\epsilon}^{\top}+\alpha{I}_{n}\otimes M_{m}$ and $\mathcal{C}_{\epsilon}+\alpha{I}_{n}\otimes M_{m}$ as its diagonal blocks, is also invertible. Therefore, $\mathcal{P}_{\epsilon}$ is invertible.
@@
The convergence behavior of the GMRES solver is closely related to Krylov subspaces. For a square matrix $\mathbf{E}\in\mathbb{R}^{k\times k}$ and a vector ${\bf y}\in\mathbb{R}^{k\times 1}$, the Krylov subspace of degree $j\geq 1$ generated by the matrix $\mathbf{E}$ and the vector ${\bf y}$ is defined as follows
@@
(see, e.g., [30]) For a non-singular $k\times k$ real linear system ${\bf Z}{\bf y}={\bf w}$, let ${\bf y}_{j}$ be the GMRES iterative solution at the $j$-th iteration $(j\geq 1)$ with ${\bf y}_{0}$ as an initial guess. Then, the $j$-th iteration solution ${\bf y}_{j}$ minimizes the residual error over the Krylov subspace $\mathcal{K}_{j}({\bf Z},{\bf r}_{0})$ with ${\bf r}_{0}={\bf w}-{\bf Z}{\bf y}_{0},$ i.e.,
@@
The following theorem establishes a connection between the convergence rate estimates of the GMRES solver applied to (17) and (16).
@@
Let $\tilde{\mathbf{x}}_{0}$ be arbitrary initial guess for (17). Let $\mathbf{x}_{0}:=\mathcal{G}^{-1}\mathcal{W}^{-\frac{1}{2}}\tilde{\mathbf{x}}_{0}$ be the initial guess for (16). Let $\mathbf{x}_{j}~{}(\tilde{\mathbf{x}}_{j},~{}{respectively})$ be the j-th $(j\geq 1)$ iteration solution derived by applying GMRES solver to (16) ((17)$,~{}{respectively})$ ) with ${\color[rgb]{0,0,0}\mathbf{x}_{0}}~{}({\color[rgb]{0,0,0}\tilde{\mathbf{x}}_{0}},~{}{respectively})$ as an initial guess. Then,
$\left\|\mathbf{r}_{j}\right\|_{2}\leq\sqrt{2}||\mathcal{W}^{-\frac{1}{2}}||_{2}||\tilde{\bf r}_{j}||_{2},\quad j=1,2,\dots,$
@@
By definition of $\tilde{\bf x}_{j}$ and $\tilde{\bf x}_{0}$, we know that $\tilde{\bf x}_{j}-\tilde{\bf x}_{0}\in\mathcal{K}_{j}(\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}},\tilde{\mathbf{r}}_{0})$ with $\tilde{\mathbf{r}}_{0}=\tilde{\bf b}-\widetilde{\mathcal{H}}_{\epsilon}^{-1}\widetilde{\mathcal{B}}\tilde{\mathbf{x}}_{0}$, meaning that there exists real numbers $\eta_{0},\eta_{1},\dots,\eta_{j-1}$ such that
@@
Denote with $\sigma({\bf C})$, the spectrum of a square matrix ${\bf C}$.
@@
Considering ~$\widetilde{\mathcal{E}}=(\widetilde{\mathcal{T}}+\alpha I_{mn})^{-1}(-\widetilde{\mathcal{T}}+\alpha I_{mn})$ with its singular value decomposition, namely, ~$\widetilde{\mathcal{E}}=\mathcal{U}\Sigma\mathcal{V}^{{\color[rgb]{0,0,0}\top}}$. Then, we can further decompose ~$\widetilde{\mathcal{H}}^{-1}\widetilde{\mathcal{B}}$ as follows:
~$\widetilde{\mathcal{H}}^{-1}\widetilde{\mathcal{B}}$ $=$ $\begin{bmatrix}I_{mn}&-\widetilde{\mathcal{E}}^{\top}\\ \widetilde{\mathcal{E}}&I_{mn}\end{bmatrix}=\begin{bmatrix}\mathcal{V}&\\ &\mathcal{U}\end{bmatrix}\begin{bmatrix}I_{mn}&-{\color[rgb]{0,0,0}\Sigma}\\ \Sigma&I_{mn}\end{bmatrix}\begin{bmatrix}\mathcal{V}^{{\color[rgb]{0,0,0}\top}}&\\ &\mathcal{U}^{{\color[rgb]{0,0,0}\top}}\end{bmatrix}.$